Normal distribution about a mean of 60 seconds and a standard deviation of 8 seconds. If 200 patient were selected at random, how many of them could be taken to the hospital in
(a) Less than 50 seconds
(b) more than 64 seconds
1
Expert's answer
2020-06-10T19:50:44-0400

The question seems to be missing the first line.


Let X = the random variable denoting the time taken to reach the hospital


We are given,


X ~ N("\mu" = 60, "\sigma"2 = 82)


Then we have,


Z = "\frac{X-\mu}{\sigma}" ~ N(0, 1), Z is the standard normal variate


(a) The probability that a patient could be taken to the hospital in less than 50 seconds


P(X < 50)


= P("\frac{X-60}{8}<\frac{50-60}{8}")


= P(Z < -1.25)


= "\Phi"(- 1.25) = 0.1056


"\therefore" If 200 patient were selected at random the number of them could be taken to the hospital in less than 50 seconds = 200 x 0.1056 = 21.12 "\approx" 21 (rounded to the nearest integer)


Answer: The number of patients could be taken to the hospital in less than 50 seconds is 21.


(b) The probability that a patient could be taken to the hospital in more than 64 seconds


P(X > 64)


= P("\frac{X-60}{8}>\frac{64-60}{8}")


= P(Z > 0.5)


= 1 - P(Z "\leq" 0.5)


= 1 - "\Phi"(0.5)


= 1 - 0.6915 = 0.3085


"\therefore" If 200 patient were selected at random the number of them could be taken to the hospital in more than 64 seconds = 200 x 0.3085 = 61.7 "\approx" 62 (rounded to the nearest integer)


Answer: The number of patients could be taken to the hospital in more than 64 seconds is 62.